During 2013, Swifty Corporation spent $174,240 in research and development costs. As a result, a new product called the New Age Piano was patented. The patent was obtained on October 1, 2013, and had a legal life of 20 years and a useful life of 10 years. Legal costs of $49,680 related to the patent were incurred as of October 1, 2013. Your answer is partially correct. Try again. Prepare all journal entries required in 2013 and 2014 as a result of the transactions above. (Round answers to 0 decimal places, e.g. 8,564. Credit account titles are automatically indented when amount is entered Do not indent manually. If no entry is required, select No Entry for the account titles and enter 0 for the amounts.) Date Account Titles and Explanation Credit arch and Development 174,2401 To record research and development expenses) To record legal expenses) ation 1,242 1,242 To record amortization expense) zation Ex Your answer is partially correct. Try again. On June 1, 2015, Swifty spent $38,880 to successfully prosecute a patent infringement suit. As a result, the estimate of useful life was extended to 12 years from June 1, 2015. Prepare all journal entries required in 2015 and 2016. (Round answers to 0 decimal places, e.g. 8,564. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter 0 for the amounts.) Date Account Titles and Explanation Debit 138,880 To record patents) 2015 zation 3,888 3,888 To record amortization expense) 2016 ation 3,888 3,888If amortization starts on Oct 1 2013, that would mean we amortize for 3 months ((49,680 / 10) * 3/12) = 1,242. Would do we not subtract this amount from 49,680 and divide by 9 to get the 2014 amortization amount? The same applies for 2015 and 2016. Could you please help me understand this? Thank you.
Answer #1

Patent is a Section 197 Intangible. You must generally amortize over 15 years the capitalized costs of "section 197 intangibles" you acquired after August 10, 1993. You must amortize these costs if you hold the section 197 intangibles in connection with your trade or business or in an activity engaged in for the production of income.

As the asset was placed in service on Oct 1, 2013, we should amortize it for 15 years or 180 months.

with this the amortization will be as follows

2013 = 49,680/15*(3/12) = $828 ( considering only 3 months for 2013)

2014 = 49,680/15 = $3,312 (we should not be taking the deduction for 9 months, since the asset was used for full year. Thus, we get to amortize for the entire year. Same follows for 2015 and 2016 and so on upto 2027.

The 9 months come into picture towards the last year i.e, 2028

In 2028, we should amortize for the first 9 months of the year. So, the amortization would be

2028 = (49,680/15)*(9/12) = $2,484.

Summing up all the years, we should get $49,680.

= $828 + (3,312*14) + $2,484 = $49,680

The first and last year [2013 and 2028 is counted as 12 months or one year ( 3 months of 2013 and 9 months of 2028)]. Rest are 14 full years.